Differentiate with respect to x


f(x) = 6x


y = 12e5x


y = ln3x


1
Expert's answer
2021-04-15T07:40:32-0400

f(x) = 6x

"\frac d{dx}(6x) = 6 \frac d{dx}(x) = 6 \cdot 1 = 6"


y = 12e5x

"\frac d{dx}(12e^{5x}) = 12\frac d{dx}(e^{5x}) = 12e^{5x}\frac d{dx}(5x) = 12e^{5x} \cdot 5 = 60e^{5x}"


y = ln3x

"\frac d{dx}(\ln3x) = \frac1{3x}\frac d{dx}(3x) = \frac1{3x}\cdot 3 = \frac1x"